# Asymptotically tight Lagrangian dual of smooth nonconvex problems via improved error bound of Shapley-Folkman Lemma

### v1

@@ 3.1 Shapley-Folkman Lemma
Shapley-Folkman Lemma is a direct consequence of Carathéodory’s theorem and there are many different ways to prove it [bi2016refined, kerdreux2023stable] and we defer our geometric proof in Lemma 4.6. Shapley-Folkman Lemma can be interpreted that the nonconvexity of the Minkowski sum of nonconvex sets, once reaching certain cap depending on the dimension, will not accumulate as the number of nonconvex sets increase. This phenomenon can be quantified by the following corollary. The proof of a slightly improved version is provided later in Lemma 4.12 under the pointedness assumption. In general, estimating the nonconvexity measures $\Phi(\cdot)$ and $\Xi(\cdot)$ is challenging. Many applications of the Shapley–Folkman Lemma therefore consider bounded nonconvex sets using the second part of Corollary 1.
@@
Let Let $A_{1},\dots,A_{k}$ be nonempty sets of $\mathbb{R}^{n}$ with $k\geq n$ such that $\Xi(A_{i})\leq\beta,\forall i\in[k]$. Then it follows that
$\Phi\left(\sum_{i\in[k]}A_{i}\right)\leq\Xi\left(\sum_{i\in[k]}A_{i}\right)\leq\sqrt{n}\beta\text{ and therefore }\Phi\left(\frac{1}{k}\sum_{i\in[k]}A_{i}\right)\leq\frac{n}{k}\beta.$
If we further assumes $\operatorname{\operatorname{R}}(A_{i})\leq\gamma,\forall i\in[k]$, then it follows that
$\Phi\left(\sum_{i\in[k]}A_{i}\right)\leq\Xi\left(\sum_{i\in[k]}A_{i}\right)\leq\sqrt{n}\gamma\text{ and therefore }\Phi\left(\frac{1}{k}\sum_{i\in[k]}A_{i}\right)\leq\frac{n}{k}\gamma.$
3.2 Duality gap of Lagrangian relaxation
In optimization theory, a key quantity associated with the Lagrangian dual that reflects the tightness of the relaxation is the duality gap:
@@ 3.2 Duality gap of Lagrangian relaxation
In this subsection, we wish to obtain a priori bound on $\Delta$ based on $\Phi(\mathcal{P})$. To achieve this, we first use the following primal characterization [boland2018combining, geoffrion2009lagrangean] of Lagrangian relaxation (bi-conjugacy in convex analysis [hiriart2004fundamentals]) to reduce $\operatorname{\operatorname{DUAL}}(\mathbf{b})$ from min-max optimization to a minimization problem in the original space. Consider
$\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})$= $\min\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}}\right\rangle$ (5)
@@
[geoffrion2009lagrangean, hiriart2004fundamentals] Under proper regularity condition, $\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})=\operatorname{\operatorname{DUAL}}(\mathbf{b})$.
Common regularity conditions required for Theorem 2 include the existence of a Slater point or the assumption that each ${\mathbf{conv}}(\mathcal{X}^{(i)})$ is a polyhedron [dey2025geoffrion, lemarechal2001geometric]. The absence of such regularity may result in cases where $\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})>\operatorname{\operatorname{DUAL}}(\mathbf{b})$ [lemarechal2001geometric]. Throughout this paper, we exclude these pathological instances and directly assume that the following equivalence holds:
@@
We finally requires two additional assumption on the geometry of $\mathcal{X}^{(i)}$.
Assumption 2.
${\overline{\mathbf{conv}}}\{\mathcal{X}^{(i)}\}={\mathbf{conv}}(\mathcal{X}^{(i)}),\forall i\in[k]$ .
Assumption 3.
[…]
We note that Assumption 2 and Assumption 3 are quite general. They are automatically satisfied when $\mathcal{X}^{(i)}$ is compact or $\mathcal{X}^{(i)}$ is an epigraph of a closed $1$-coercive function. Even if each set $\mathcal{X}^{(i)}$ is closed, it is still possible that its convex hull ${\mathbf{conv}}(\mathcal{X}^{(i)})$ is not closed. Therefore, the closed convex hull operator is required in (5) in general. Since our goal is to compare (5) with (1), additional structures or Assumption 2 are expected. On the other hand, Assumption 3 guarantees that each ${\mathbf{conv}}(\mathcal{X}^{(i)})$ has at least one extreme point, and moreover any such extreme point must lie in the original set $\mathcal{X}^{(i)}$. The pointedness assumption is subsequently used in Lemma 4.6 to describe how the Minkowski sum of pointed closed convex sets decomposes in terms of their facial structure. This decomposition provides a geometric proof of the Shapley–Folkman Lemma. We remark that many results in this paper continue to hold—after modest modifications—even without Lemma 4.6 and Assumption 3. Nevertheless, we include Lemma 4.6 because it gives a clean geometric characterization of Minkowski sums of pointed closed convex sets and conveys the central theme of this paper: the local geometry around extreme points impacts the global behavior of Minkowski sums.
As we mentioned earlier, the tightness of Lagrangian dual of (1) directly relies on the convexity of $\mathcal{P}=\sum_{i=1}^{k}\mathcal{P}^{(i)}$ and we close this section by quantifying $\Delta$ in terms of $\Phi(\mathcal{P})$.
@@
Let $\mathcal{E}:=\Phi\left(\mathcal{P}\right)$. Under Assumption 1 and Assumption 2, it then follows that
@@
Under the Assumption 1, we know that $\operatorname{\operatorname{DUAL}}(\mathbf{b})=\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})$.
Let $\mathbf{x}_{\textbf{L}}$ be the optimal solution of (5). By weak duality, we know that
$\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})=\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{\textbf{L}}}\right\rangle=\operatorname{\operatorname{DUAL}}(\mathbf{b})\leq\operatorname{\operatorname{OPT}}(\mathbf{b}).$ (6)
Let $\mathbf{p}^{(i)}_{\textbf{L}}:=\begin{bmatrix}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{\textbf{L}}}\right\rangle\\ B^{(i)}\mathbf{x}^{(i)}_{\textbf{L}}\end{bmatrix}$. By Claim 3 and Assumption 2, $\mathbf{p}^{(i)}\in{\mathbf{conv}}(\mathcal{P}^{(i)})$. By Claim 2, $\sum_{i=1}^{k}\mathbf{p}^{(i)}\in\sum_{i=1}^{k}{\mathbf{conv}}(\mathcal{P}^{(i)})={\mathbf{conv}}(\mathcal{P})$. By the definition of $\Phi\left(\mathcal{P}\right)$, there exists some $\mathbf{x}_{*}\in\mathcal{X}^{(i)}$ and $\mathbf{p}^{(i)}_{*}:=\begin{bmatrix}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{*}}\right\rangle\\ B^{(i)}\mathbf{x}^{(i)}_{*}\end{bmatrix}$ such that
$\left\lVert\left(\sum_{i=1}^{k}\mathbf{p}^{(i)}_{\textbf{L}}\right)-\left(\sum_{i=1}^{k}\mathbf{p}^{(i)}_{\textbf{*}}\right)\right\rVert_{2}\leq\mathcal{E}.$
@@
$\operatorname{\operatorname{OPT}}(\mathbf{b}+\mathcal{E}\textbf{1})\leq\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{*}}\right\rangle\leq\left(\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{\textbf{L}}}\right\rangle\right)$$\mathcal{E}$.
Combining with (6), this yields that
$\Delta=\operatorname{\operatorname{OPT}}(\mathbf{b})-\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})\leq\mathcal{E}+\operatorname{\operatorname{OPT}}(\mathbf{b})-\operatorname{\operatorname{OPT}}(\mathbf{b}+\mathcal{E}\textbf{1}).$
@@ 4.1 Geometric proof of Shapley-Folkman Lemma
In this section, we give a proof of Shapley-Folkman Lemma that utilized the facial structure of the convex hull of each nonconvex sets.
@@
Given a convex set $C$, we call a close convex set $F\subseteq C$ is a face of $C$ if for any line segment $a,b]\subseteq C$ such that $(a,b)\cap F\neq\emptyset$, we have $a,b]\in F$. If there exists a supporting hyperplane $H$ of $C$ such that $F=C\cap H$, then we call $F$ an exposed face. The dimension of face $F$ denoted by $\dim(F)$ is the dimension of its affine hull.
@@
Proof 4.7.
For any $\mathbf{y}\in{\mathbf{conv}}\left(\sum_{i=1}^{k}A_{i}\right)$, consider $\mathbf{y}=\sum_{i=1}^{k}\mathbf{y}_{i}$ where each $\mathbf{y}_{i}$ lies in the relative interior some face $F_{i}$ of ${\mathbf{conv}}(A_{i})$. If $\sum_{i=1}^{k}\dim(F_{i})>n$, we show how to find a different $\mathbf{y}_{i}^{\prime}$ from ${\mathbf{conv}}(A_{i})$ so that the total dimension $\sum_{i=1}^{k}\dim(F_{i})$ is decrease. For each $F_{i}$, we can find a set $V_{F_{i}}$ of $\dim(F_{i})$ linear independent vectors so that $\mathbf{y}_{i}\pm\epsilon\mathbf{v}\in F_{i}$ for all $\mathbf{v}\in V_{F_{i}}$ and small enough $\epsilon>0$. Now extending $V_{F_{i}}$ in the space of $A_{1}\times A_{2}\times\cdots A_{k}$, we construct
$V=\left\{\tilde{{\mathbf{v}}}:=\begin{bmatrix}\textbf{0}\\ \mathbf{v}\\ \textbf{0}\\ \end{bmatrix}\in\mathbb{R}^{(i-1)n}\times\mathbb{R}^{n}\times\mathbb{R}^{(k-i)n},\forall\mathbf{v}\in V_{F_{i}},\forall i\in[k]\right\}.$
Clearly, $V$ includes $\sum_{i=1}^{k}\dim(F_{i})$ linearly independent vectors. If $\sum_{i=1}^{k}\dim(F_{i})>n$, there exists $\mathbf{d}=[\mathbf{d}_{1},\mathbf{d}_{2},\cdots,\mathbf{d}_{k}]\in\text{span}\{V\}$ that
$\mathbf{d}_{1}+\mathbf{d}_{2}+\cdots+\mathbf{d}_{k}=\textbf{0}$
Let $\mathbf{y}_{1}^{\prime},\mathbf{y}_{2}^{\prime},\cdots,\mathbf{y}_{k}^{\prime}]=[\mathbf{y}_{1},\mathbf{y}_{2},\cdots,\mathbf{y}_{k}]+\lambda\mathbf{d}$. Since each ${\mathbf{conv}}(A_{i})$ is pointed, there exists some choice of $\lambda,\epsilon$ such that at least one $\mathbf{y}_{i}^{\prime}$ reaches the boundary of $F_{i}$ while keeping other $\mathbf{y}_{j}^{\prime}$ in $F_{j}$. In this case, the total dimension $\sum_{i=1}^{k}\dim(F_{i})$ is decreased and we still ensure that $\mathbf{y}=\sum_{i=1}^{k}\mathbf{y}_{i}^{\prime}$ for some $\mathbf{y}^{\prime}_{i}\in{\mathbf{conv}}(A_{i})$.
Since $\dim(F_{i})\in\{0,1,\dots,n\}$ and $\sum_{i=1}^{k}\dim(F_{i})\leq n$ implies that at least knk-n faces have dimension 0 and are extreme points which lies in $A_{i}$. Let $\mathcal{I}=\{i\in[k]:\dim(F^{(i)})>0\}$. It is straightforward to check that
@@
and therefore Lemma 4.6 recovers the class Shapley-Folkman Lemma (Theorem 1) under the pointedness condition.
Moreover, one can further use Lemma 4.6 to strength the refined Shapley-Folkman Lemma recently proposed in [bi2016refined] where the notion of $r$-th convex hull in [bi2016refined] corresponds to dimension of face $F_{i}$ in Lemma 4.6 but we additionally requires that the points used to construct $r$-th convex hull lie in a $r$-dimension face $F_{i}$.

### v2

@@ 2.3 Shapley-Folkman Lemma
Shapley-Folkman Lemma is a direct consequence of Carathéodory’s theorem and there are many different ways to prove it [13, 14] and we defer our geometric proof in Lemma 4.7, presented in Section 4.
Shapley-Folkman Lemma can be interpreted that the nonconvexity of the Minkowski sum of nonconvex sets, once reaching certain cap depending on the dimension, will not accumulate as the number of nonconvex sets increase. This phenomenon can be quantified by the following corollary. In general, estimating the nonconvexity measures $\Phi(\cdot)$ and $\Xi(\cdot)$ is challenging. Many applications of the Shapley–Folkman Lemma therefore consider bounded nonconvex sets using the second part of Corollary 1.
@@
Let $A_{1},\dots,A_{k}$ be nonempty sets of $\mathbb{R}^{n}$ with $k\geq n$ such that $\Xi(A_{i})\leq\beta,\forall i\in[k]$. Then it follows that
$\Phi\left(\sum_{i\in[k]}A_{i}\right)\leq\Xi\left(\sum_{i\in[k]}A_{i}\right)\leq\sqrt{n}\beta\text{ and therefore }\Phi\left(\frac{1}{k}\sum_{i\in[k]}A_{i}\right)\leq\frac{\sqrt{n}}{k}\beta.$
If we further assume $\operatorname{\operatorname{R}}(A_{i})\leq\gamma,\forall i\in[k]$, then it follows that
$\Phi\left(\sum_{i\in[k]}A_{i}\right)\leq\Xi\left(\sum_{i\in[k]}A_{i}\right)\leq\sqrt{n}\gamma\text{ and therefore }\Phi\left(\frac{1}{k}\sum_{i\in[k]}A_{i}\right)\leq\frac{\sqrt{n}}{k}\gamma.$
3 Shapley-Folkman Lemma and duality gap of Lagrangian relaxation
A key quantity associated with the Lagrangian dual that reflects the tightness of the relaxation is the duality gap (also called the additive gap) which is defined as:
@@ 3 Shapley-Folkman Lemma and duality gap of Lagrangian relaxation
In this section, we present a result showing how to upper bound $\Delta$ based on upper bounds on $\Phi(\mathcal{P})$.
To achieve this, we first use the following primal characterization [19, 20] of Lagrangian dual (bi-conjugacy in convex analysis [21]) to reduce $\operatorname{\operatorname{DUAL}}(\mathbf{b})$ from min-max optimization to a minimization problem in the original space. Consider
$\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})$= $\inf\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}}\right\rangle$ (5)
@@
[20, 21] Under proper regularity condition, $\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})=\operatorname{\operatorname{DUAL}}(\mathbf{b})$.
Common regularity conditions required for Theorem 2 include the existence of a Slater point or the assumption that each ${\mathbf{conv}}(\mathcal{X}^{(i)})$ is a polyhedron [22, 23]. The absence of such regularity may result in cases where $\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})>\operatorname{\operatorname{DUAL}}(\mathbf{b})$ [23]. Throughout this paper, we exclude these pathological instances and directly assume that the following equivalence holds:
@@
We finally present the main result of this Section showing that the tightness of Lagrangian dual of (1) directly relies on the convexity of $\mathcal{P}=\sum_{i=1}^{k}\mathcal{P}^{(i)}$ by quantifying $\Delta$ in terms of $\Phi(\mathcal{P})$. This theorem is similar to the gap formula in [24].
@@
Let $\mathcal{E}:=\Phi\left(\mathcal{P}\right)$. Under Assumption 1 and ${\mathbf{conv}}(\mathcal{X}^{(i)})$ being closed for all $i\in[k]$, it follows that
@@
Under Assumption 1, we know that $\operatorname{\operatorname{DUAL}}(\mathbf{b})=\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})$. Let $\mathbf{x}_{\textbf{L}}$ be the $\epsilon$-optimal solution of (5). By weak duality, we know that
$\operatorname{\operatorname{OPT}}_{\textbf{L}}(\mathbf{b})+\epsilon=\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{\textbf{L}}}\right\rangle=\operatorname{\operatorname{DUAL}}(\mathbf{b})+\epsilon\leq\operatorname{\operatorname{OPT}}(\mathbf{b})+\epsilon.$ (6)
Let $\mathbf{p}^{(i)}_{\textbf{L}}:=\begin{bmatrix}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{\textbf{L}}}\right\rangle\\ B^{(i)}\mathbf{x}^{(i)}_{\textbf{L}}\end{bmatrix}\in{\mathbf{conv}}\left(\mathcal{P}^{(i)}\right)$ where the last containment is due to the fact that ${\mathbf{conv}}(\mathcal{X}^{(i)})$ is closed (which implies $\mathbf{x}^{(i)}_{\textbf{L}}\in{\mathbf{conv}}(\mathcal{X}^{(i)})$) and using Lemma 3. By Lemma 2, $\sum_{i=1}^{k}\mathbf{p}^{(i)}\in\sum_{i=1}^{k}{\mathbf{conv}}\left(\mathcal{P}^{(i)}\right)={\mathbf{conv}}(\mathcal{P})$. By the definition of $\Phi\left(\mathcal{P}\right)$, there exists some $\mathbf{x}_{*}\in\mathcal{X}^{(i)}$ and $\mathbf{p}^{(i)}_{*}:=\begin{bmatrix}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{*}}\right\rangle\\ B^{(i)}\mathbf{x}^{(i)}_{*}\end{bmatrix}$ such that
$\left\lVert\left(\sum_{i=1}^{k}\mathbf{p}^{(i)}_{\textbf{L}}\right)-\left(\sum_{i=1}^{k}\mathbf{p}^{(i)}_{\textbf{*}}\right)\right\rVert_{2}\leq\mathcal{E}.$ (7)
@@
$\operatorname{\operatorname{OPT}}(\mathbf{b}+\mathcal{E}\textbf{1})\leq\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{*}}\right\rangle\leq\left(\sum_{i=1}^{k}\left\langle{\mathbf{c}^{(i)}},{\mathbf{x}^{(i)}_{\textbf{L}}}\right\rangle\right)$$\mathcal{E}$,
where the last inequality follows from (7). Now, using (6) with the above inequality, we obtain that
$\operatorname{\operatorname{OPT}}(\mathbf{b}+\mathcal{E}\textbf{1})-\mathcal{E}\leq\operatorname{\operatorname{DUAL}}(\mathbf{b})+\epsilon$
$\Rightarrow$ $\Delta\leq\mathcal{E}+\operatorname{\operatorname{OPT}}(\mathbf{b})+\epsilon-\operatorname{\operatorname{OPT}}(\mathbf{b}+\mathcal{E}\textbf{1}),$
where the second inequality is obtained by subtracting $\operatorname{\operatorname{OPT}}(\mathbf{b})$ on both sides. Letting $\epsilon\to 0$ yields the desired statement.
@@ 4 Error bounds via Shapley-Folkman Lemma
In this section (and for the rest of the paper), we make an additional assumption:
Assumption 2.
Let $A_{i}\subseteq\mathbb{R}^{n}$, $i\in[k]$ be the sets that are the summands in the Shapley–Folkman lemma. For all $i\in[k]$, we assume that ${\mathbf{conv}}(A_{i})$ is closed and pointed.
Assumption 2 guarantees that each ${\mathbf{conv}}(A_{i})$ has at least one extreme point. Moreover, since ${\mathbf{conv}}(A_{i})$ is closed, any such extreme point must lie in the original set $A_{i}$. The pointedness assumption is subsequently used in Lemma 4.7 in the next section to describe how the Minkowski sum of pointed closed convex sets decomposes in terms of their facial structure. This decomposition provides a geometric proof of the Shapley–Folkman Lemma.
@@ 4.1 Geometric proof of Shapley-Folkman Lemma
In this section, we give a proof of Shapley-Folkman Lemma that utilizes the facial structure of the convex hull of each nonconvex set.
For completeness, we begin by presenting the standard definition of a face of a convex set and related notions.
@@
Given a closed convex set $C$, we call a closed convex set $F\subseteq C$ is a face of $C$ if for any line segment $a,b]\subseteq C$ such that $(a,b)\cap F\neq\emptyset$, we have $a,b]\in F$. The dimension of face $F$ denoted by $\dim(F)$ is the dimension of its affine hull.
Note that a convex set $C$ is a face of itself. Moreover a face of a face of $C$ is also a face of $C$. Finally, faces of closed convex sets are closed convex sets [21].
We begin some simple observations from linear algebra and convex analysis. The first claim directly follows from a rank-nullity argument.
Claim 1.
Let $L_{i}\in\mathbb{R}^{n}$, $i\in[k]$ be linear subspaces, such that $\sum_{i=1}^{k}\dim(L_{i})>n$. Then there exists $d_{i}\in L_{i}$ for $i\in[k]$, with not all $d_{i}$s zero, such that $\sum_{i=1}^{k}d_{i}=0.$
Claim 2.
Let $C\subseteq\mathbb{R}^{n}$ be a closed convex set, and let $F$ be a nonempty face of $C$. If $x\in\textup{relbd}(F)$, then $x$ is contained in a face of $C$ of strictly smaller dimension than $F$.
Proof 4.6.
Since $x\in\textup{relbd}(F)$, there exists a hyperplane $H$ in the affine hull of $F$ such that (i) $F$ is contained in one of the half-spaces defined by $H$, (ii) $x\in H$, and (iii) $F\not\subseteq H$. Therefore, by definition of a face, $F\cap H$ is a face of $F$, and thus a face of $C$. Moreover, $(iii)$ implies that the dimension of $F$ is at least 1 more than the dimension of $F\cap H$.
@@
Proof 4.8.
[…]
Proof 4.9.
of Shapley–Folkman Lemma using Lemma 4.7 For every $\mathbf{y}\in{\mathbf{conv}}\left(\sum_{i=1}^{k}A_{i}\right)$, Lemma 4.7 implies that there exist faces $F_{i}$ of ${\mathbf{conv}}(A_{i})$ for $i\in[k]$, such that $\mathbf{y}\in\sum_{i=1}^{k}F_{i}\text{ and }\sum_{i=1}^{k}\dim(F_{i})\leq n.$ Since $\dim(F_{i})\in\{0,1,\dots,n\}$ and $\sum_{i=1}^{k}\dim(F_{i})\leq n$, it implies that at least knk-n faces have dimension 0. These dimension-0 faces are extreme points of corresponding ${\mathbf{conv}}(A_{i})$s, therefore must lie in $A_{i}$. Let $\mathcal{I}=\{i\in[k]:\dim(F_{i})>0\}$ where from the above we have that $|\mathcal{I}|\leq n$. Moreover, we have,
@@
We note here that Lemma 4.7 can be used to strengthen the refined Shapley-Folkman Lemma recently proposed in [13].